OpenStudy (anonymous):

A truck transports 7 cars to an auto dealership. Each car has a gasoline tank capacity of 64 liters. What is the total gasoline tank capacity of all of the cars in kiloliters?

OpenStudy (anonymous):

@ash2326 help plz

OpenStudy (hoblos):

64*7 = 448 liters 1kiloliter = 1000liter 448 liters= 448/1000= 0.448 kiloliters

OpenStudy (anonymous):

o thx so much

OpenStudy (ash2326):

Truck transports 7 cars to an auto dealer. Each car has a gasoline tank capacity of 64 liters. \[Total\ gas\ capacity= No.\ of\ cars\ \times\ capacity\ of\ each\ car\] Can you do now?

OpenStudy (ash2326):

For kilo liters , divide the answer obtained by 1000

OpenStudy (anonymous):

ok thx